# Quaternion Multi-focus Color Image Fusion

### v1

@@ IV Experiments
To evaluate the fusion performance of our framework, we conduct experiments on three public datasets. Section A presents experimental settings like competing methods, datasets, and evaluation metrics. Sections B and C discuss the convergence and optimal parameters of the QSD model. To demonstrate the superiority of our method, Sections D and E present experimental results compared with other state-of-the-art approaches. Section F presents ablation studies on the mffw dataset [30] to verify the effectiveness of our QSD model, QBDF and QSSR strategies.
@@ IV-A Experimental settings
In our experiments, the patch size is set as $\sqrt{d}\times\sqrt{d}$ where $\sqrt{d}=8$ by default in the QSD model and base-scale fusion process. In the detail-scale fusion and the QSSR process, $\sqrt{d}$ is set to $5\times 10^{-3}MN$ [18], where $M$ and $N$ denote the height and width of the input image. The stopping criterion of QSD model is $51e-$ 5 in Algorithm 1. All experiments are performed using Matlab 2016 on a workstation with a 2.90GHz Intel Core CPU and 16GB memory.
Comparison methods. Nine representative multi-focus image fusion algorithms are selected as the comparison methods. They include sparse representation-based methods (SR) [10] as the baseline method, Guide-filter-based method (GFDF) [11], Multi-scale gradients-based method (MGIMF) [16], GAN-based method (MFIF-GAN) [31], Higher-order SVD-based method (QHOSVD) [19], unsupervised zero-shot based method (ZMFF) [32], Small-area-aware method (SAMF) [4], End-to-End based network (DBMFIF) [7], General-Image-Fusion-based (TCMOA) [33]. SR belongs to channel-wise processing-based MCIF methods. GFDF, MGIMF, and SAMF are grayscale conversion-based MCIF methods. QHOSVD is a quaternion domain-based MCIF method. MFIF-GAN, ZMFF, DBMFIF, and TCMOA are deep-learning-based MCIF methods.
@@
Metrics. Following [2], six objective metrics are adopted to evaluate the performance of two-color image fusion quantitatively. They are the normalized mutual information $Q_{MI}$, gradient-based metric $Q_{G}$, phase congruency-based metric $Q_{P}$, structural similarity-based metric $Q_{E}$, $Q_{Y}$ and human perception-based metric $Q_{CB}$. $Q_{MI}$ quantifies the intensity-based mutual information transferred from source images to the fused output. $Q_{G}$ and $Q_{P}$ assess the preservation of gradient information and salient feature details. $Q_{Y}$ evaluates the structural preservation of the fused image in terms of luminance, contrast, and texture patterns while $Q_{E}$ extends $Q_{Y}$ by considering edge consistency in the fused result. $Q_{CB}$ measures the preservation of contrast variations from source images in the fused image.
IV-B Convergence Analysis
@@
The convergence analysis of the QSD model is provided in this subsection. Fig. 4 displays the evolution curves of the relative difference versus iterations of QSD model on the samples of lytro [34], mffw [30] and MFI-WHU datasets [35]. The relative difference is calculated by the maximum value between $\|\dot{\bm{\mathrm{Z}}}{{}^{t+1}}-\dot{\bm{\mathrm{Z}}}{{}^{t}}\|{{}_{\infty}}$ and $\|\dot{\bm{\mathrm{D}}}{{}^{t+1}}-\dot{\bm{\mathrm{D}}}{{}^{t}}\|{{}_{\infty}}$. $\dot{\bm{\mathrm{Z}}}{{}^{t+1}}$ and $\dot{\bm{\mathrm{Z}}}{{}^{t}}$ denote the coefficient matrices at successive $t^{th}$ and ${(t+1)}^{th}$ iterations respectively. $\dot{\bm{\mathrm{D}}}{{}^{t+1}}$ and $\dot{\bm{\mathrm{D}}}{{}^{t}}$ denote the detail layers at successive $t^{th}$ and ${(t+1)}^{th}$ iterations respectively. The $\infty$ norm means the largest magnitude among each element of a vector. To improve the QSD speed, $\dot{\bm{\mathrm{B}}}{{}^{0}}$ is initialized as $\frac{\mathcal{F}(\dot{\bm{\mathrm{I}}})}{\mathcal{F}(\nabla_{1}^{T}\nabla_{1}+\nabla_{2}^{T}\nabla_{2})+1}$ and $\dot{\bm{\mathrm{D}}}{{}^{0}}$ is set to $\dot{\bm{\mathrm{I}}}-\dot{\bm{\mathrm{B}}}{{}^{0}}$ As displayed in Fig. 4, QSD is convergent and its total number of iterations is 20. This can sufficiently guarantee convergence.
@@
Figure 4: Convergence analysis of QSD model on lytro, mffw and MFI-WHU datasets.
IV-C Parameter Discussions
Here, we discuss how three regularization parameters $\alpha$, $\beta$, and $\lambda$ affect the fusion performance of the QSD model. First, we show the visual effects of various values of $\alpha$, $\beta$ and $\lambda$. Then we comprehensively analyze and determine parameter settings.
@@ IV-C Parameter Discussions
Figure 5: The fusion performance with different parameters on the lytro and the mffw datasets. (a) $Q_{E}+Q_{CB}$ results when parameters $\alpha$ and $\beta$ vary on the mffw dataset. (b) The $Q_{MI}$ value with respect to parameter $\lambda$ on both lytro and mffw datasets.
@@
In our QSD model, parameter $\alpha$ controls the base-scale layer while parameter $\beta$ handles the detail-scale layer. The final fused image not only preserves the most detailed parts of the inputs but also maintains the consistency of the background.
@@
To handle the three regularization parameters in our QSD model, we fix parameter $\lambda=0.05$ and adjust the others each time. $\alpha$ and $\beta$ were sequentially set values of [0.5, 1, 1.5, 2, 2.5, 3]. Then $\alpha$ and $\beta$ were set to 1.5 and 2 and the variation values of $\lambda$ were sequentially set values of [0.05, 0.06, 0.07, 0.08, 0.09, 0.1]. To effectively estimate the fusion performance, the combination index $Q_{E}+Q_{CB}$ approximates the fused images in terms of detail and base parts. Since $\lambda$ is not remarkably sensitive to the final fusion quality, index $Q_{MI}$ is used to quantitatively analyze the fused results.
To quantitatively analyze the sensitivity of QSD versus these parameters, Fig. 5 displays the $Q_{E}+Q_{CB}$ results of QSD with the samples of mffw dataset. In Fig. 5 (a), $\alpha$ and $\beta$ are set to 1.5 and 2 on mffw dataset and $\beta$ are set to 0.5 on lytro and MFI-WHU datasets. $\lambda=5e-2$ is set for all datasets in Fig. 5 (b).
@@
TABLE II: Performance comparison of multi-focus image fusion methods in two-color image fusion.
@@
Ours 1.1656 0.7613 0.8486 0.8813 0.9896 0.8056
mffw [30] SR [10] 0.9181 0.6278 0.5500 0.8078 0.7844 0.6503
GFDF [11] 1.0878 0.7323 0.7630 0.8242 0.9522 0.7426
@@
IV-D Two color image fusion
This subsection discusses the quantitative and qualitative results to assess fusion quality of two-color image fusion.
Quantitative evaluation. Table II provides the average performance of various competing algorithms on three public datasets. Six evaluation metrics are employed to assess the quality of the fused images, with higher values indicating better performance. In this paper, the highest evaluation scores are highlighted in bold.
On the lytro dataset [34], our framework achieves the highest metric values compared to the other competing methods across gradient-based, structural similarity-based, and human-perception-based evaluation metrics. This improvement is attributed to accurate focus detection capability of our method. However, our framework is slightly worse than SAMF on the metric of $Q_{MI}$. This is because our patch-wise fusion strategy tends to preserve less image intensity information during the quaternion fusion and refinement process compared to the pixel-wise fusion strategy.
Since the images on the mffw dataset are affected by severe defocus spread effects and contain more challenging low-detail regions [30], all competing methods perform worse on this dataset compared to lytro and MFI-WHU datasets. Our method achieves the highest fusion performance on $Q_{G}$, and $Q_{E}$ metrics on the mffw dataset, highlighting its superior capability in focus detection and detail preservation. In the $Q_{MI}$ results, our framework slightly outperforms GFDF thanks to the effective artifact removal of our QSSR strategy. On the metrics of $Q_{Y}$ and $Q_{CB}$, our framework is worse only than MFIF-GAN. This demonstrates the advantages of fusion and refinement provided by the GAN-based MCIF algorithm in mitigating defocus spread effects [31].
@@ IV-D Two color image fusion
On the MFI-WHU dataset, our method achieves the best fusion performance in $Q_{G}$ and $Q_{E}$ metrics, underscoring its effectiveness in detecting focus regions from small areas, particularly when compared to SAMF. However, in the $Q_{Y}$ metric, our method performs slightly worse than GFDF and MFIF-GAN since both methods exhibit strong focus detection capabilities.
@@
Extensive experiments on various datasets demonstrate our framework significantly outperforms QHOSVD in terms of detail preservation and artifact removal under real-world challenging scenarios.
Visual evaluation. To evaluate the effectiveness of detail preservation and artifact removal, our method is compared with nine state-of-the-art methods. The red and yellow bounding boxes highlight areas with quality degradation. They are enlarged to better illustrate differences in the focused and defocused regions across the input images and output fused results.
The first two rows of Fig. 6 present the fused results of all competing methods on the lytro dataset. The red bounding boxes reveal that the results of SR, MGIMF, SAMF, and QHOSVD exhibit significant blurring along the edges. MGIMF shows spatial artifacts in the regions within the yellow bounding box. This is because the suboptimal decision maps are generated by multi-scale hand-designed filters. Our framework generates a pleasant fused result with clear objects and without any artifacts.
The third and the fourth rows of Fig. 6 illustrate the fused results of all competing methods on the mffw dataset. The red bounding boxes indicate that the outputs of SR, GFDF, and QHOSVD contain numerous blurred pixels. Similarly, the yellow bounding box highlights pseudo-edges and ghosts in the boundary regions of the MGIMF result. Furthermore, the outputs of SR, GFDF, MFIF-GAN, MGIMF, and QHOSVD suffer from varying degrees of color distortion and spatial artifacts in low-gradient regions. Our framework gives the most satisfactory fused result because our framework sufficiently handles the low-grafient regions and prevents visual artifacts for uncertain boundary areas.
The last two rows of Fig. 6 show the fused images produced by all comparative methods on the MFI-WHU dataset. The input images contain small and near blurred areas. This makes it difficult to estimate the focus levels and accurately identify focused regions. As can be seen in the red and yellow bounding boxes of the results, it is evident that SR, GFDF, MFIF-GAN, MGIMF, QHOSVD, SAMF and TCMOA produce unclear regions such as electric wires and poles. In contrast, our method demonstrates superior accuracy in detecting and preserving these focused regions.
IV-E Multiple color image fusion
To verify the superiority of our method in multiple-color image fusion, we conduct the experiment on the mffw3 dataset.
@@ IV-E Multiple color image fusion
Figure 7: Three partially focused color image fusion on $mffw3$ dataset. The source 1 image is a near-focus image. The source 2 data is a middle-focused image. The source 3 data is a far-focused image.
Qualitative evaluation. Since TCMOA [33] does not provide a pre-trained model, only nine methods are included in this comparison. The deep learning-based methods fuse the multiple color images according to [36].
As displayed in Fig. 7, the experimental results show that these comparison methods perform poorly in the multiple-color image fusion process. GFDF, MGIMF, ZMFF and SAMF contain noticeable visual artifacts in the background areas. In the red bounding boxes of the QHOSVD and MGIMF results, significant blurring is observed, particularly in the areas of the wings. Similarly, in the yellow bounding box of the SAMF result, the fused image is obstructed by defocused pixels within the patterns. SR shows obvious artifacts and color distortion. This may be because SR fails to provide accurate focus level for focus detection and it is highly sensitive to large misregistration and color differences in the source images. Fig. 7 indicates the superiority of our framework in terms of focus detection and artifact removal in foreground, background and boundary regions.
IV-F Ablation study
Effectiveness of QSD. To verify the advantages of the QSD model, we perform the ablation experiments on the mffw dataset (the first 10 pairs of samples) with two input color images. Different settings of parameters in Eq. (3) are provided, we investigate how the different terms of QSD influence the overall fusion performance. Specifically, three different settings of $\alpha=0$, $\beta=0$, $\lambda=0$ and $\alpha\neq 0$, $\beta\neq 0$, $\lambda\neq 0$ are particularly considered in this experiment. Table III gives the quantitative fusion results with different settings of QSD. As can be observed, the different settings of terms can influence the fusion performance at different degrees. The QSD model provides the best quantitative results. Therefore, this experiment further verifies the effectiveness of the $\alpha$ term, $\beta$ term and $\lambda$ term in Eq. (3).
TABLE III: Quantitative evaluation on the different settings of the QSD model, the QBDF and the QSSR strategies.

### v2

@@ IV Experiments
To evaluate the fusion performance of our QMCIF framework, we conduct experiments on three public datasets. Section A presents experimental settings like competing methods, datasets, and evaluation metrics. Section B properties of the quaternion focal element decomposition (QFED) module that is a key component of the quaternion consistency-aware focus detection (QCAFD) method. To demonstrate the superiority of our QMCIF framework, Section C presents experimental results compared with state-of-the-art approaches. Section D presents ablation studies on the mffw dataset [29] to verify the effectiveness of our QCAFD module, QBDF and QSSR strategies.
@@ IV-A Experimental settings
In our experiments, the patch size is set as $\sqrt{d}\times\sqrt{d}$ where $\sqrt{d}=8$ by default in the QCAFD method and base-scale fusion process. In the detail-scale fusion and QSSR process, $\sqrt{d}$ is set to $5\times 10^{-5}MN$ [30], where $M$ and $N$ denote the height and width of the input image. The stopping criterion of QFED is 1e51e-5 in Algorithm 1. All experiments are performed using Matlab 2016 on a workstation with a 2.90GHz Intel Core CPU and 16GB memory.
Competing methods. Eleven representative multi-focus image fusion algorithms are selected as the comparison methods. They include a sparse coding-based (SR) method [14], a convolutional sparse coding-bsed (CSR) method [31], a multi-scale gradients-based (MGIMF) method [15], GAN-based (MFIF-GAN) method [6], a higher-order SVD-based (QHOSVD) method [21], unsupervised zero-shot based method (ZMFF) [32], a small-area-aware (SAMF) method [4], End-to-End based network (DBMFIF) [7], General-Image-Fusion-based (TCMOA) [33], a Swintransformer-based network (SwinMFF) [8] and a convolutional sparse representation-based unfolding network (MCCSR-Net) [34]. SR and CSR are channel-wise processing-based methods. MGIMF and SAMF are grayscale conversion-based methods. QHOSVD is a quaternion domain-based method. MFIF-GAN, ZMFF, DBMFIF, TCMOA, SwinMFF and MCCSR-Net are deep-learning-based methods. TCMOA and SwinMFF are cross-channel attention-based methods.
@@
Metrics. Following [1], six objective metrics are adopted to evaluate the performance of color image fusion quantitatively. They are the normalized mutual information $Q_{MI}$, gradient-based metric $Q_{G}$, phase congruency-based metric $Q_{P}$, structural similarity-based metric $Q_{E}$, $Q_{Y}$ and human perception-based metric $Q_{CB}$. $Q_{MI}$ quantifies the intensity-based mutual information transferred from source images to the fused output. $Q_{G}$ and $Q_{P}$ assess the preservation of gradient information and salient feature details. $Q_{Y}$ evaluates the structural preservation of the fused image in terms of luminance, contrast, and texture patterns while $Q_{E}$ extends $Q_{Y}$ by considering edge consistency in the fused result. $Q_{CB}$ measures the preservation of contrast variations from source images in the fused image.
IV-B Analysis of QFED
This subsection analyzes our quaternion focal element decomposition (QFED) model that is a key component of QCAFD. We aim to validate QFED’s convergence, parameter sensitivity, and the impact of its each term on the whole fusion performance.
@@ IV-B Analysis of QFED
Convergence Analysis. Fig. 4 displays the evolution curves of the relative difference versus iterations of QFED on the samples of lytro [35], mffw [29] and MFI-WHU datasets [36]. The relative difference is defined by the maximum value between $\|\dot{\bm{\mathrm{Z}}}{{}^{t+1}}-\dot{\bm{\mathrm{Z}}}{{}^{t}}\|{{}_{\infty}}$ and $\|\dot{\bm{\mathrm{D}}}{{}^{t+1}}-\dot{\bm{\mathrm{D}}}{{}^{t}}\|{{}_{\infty}}$. $\dot{\bm{\mathrm{Z}}}{{}^{t+1}}$ and $\dot{\bm{\mathrm{Z}}}{{}^{t}}$ denote the coefficient matrices at successive $t^{th}$ and ${(t+1)}^{th}$ iterations respectively. $\dot{\bm{\mathrm{D}}}{{}^{t+1}}$ and $\dot{\bm{\mathrm{D}}}{{}^{t}}$ denote the detail layers at successive $t^{th}$ and ${(t+1)}^{th}$ iterations respectively. The $\infty$ norm means the largest magnitude in each element of a vector. To improve the QFED speed, $\dot{\bm{\mathrm{B}}}{{}^{0}}$ is initialized as $\frac{\mathcal{F}(\dot{\bm{\mathrm{I}}})}{\mathcal{F}(\nabla_{1}^{T}\nabla_{1}+\nabla_{2}^{T}\nabla_{2})+1}$ and $\dot{\bm{\mathrm{D}}}{{}^{0}}$ is set to $\dot{\bm{\mathrm{I}}}-\dot{\bm{\mathrm{B}}}{{}^{0}}$ . As displayed in Fig. 4, QFED is converges within 20 iterations. This demonstrates the computational efficiency and stability of the iterative solution of QFED.
@@
Figure 4: Convergence analysis of QFED model on lytro, mffw and MFI-WHU datasets.
Parameter Sensitivity. We analyze the influence of regularization parameters $\alpha$, $\beta$, and $\lambda$ in the QFED model in Eq. (3).
@@
Figure 5: Fusion performance with different parameters on the lytro and the mffw datasets. (a) $Q_{E}+Q_{CB}$ results when parameters $\alpha$ and $\beta$ vary on the mffw dataset. (b) The $Q_{MI}$ value with respect to parameter $\lambda$ on both lytro and mffw datasets.
@@
In QFED, parameter $\alpha$ controls the base-scale layer while parameter $\beta$ handles the detail-scale layer. The final fused image not only preserves the most detailed regions of the inputs but also maintains the consistency of the background.
@@
To handle the three regularization parameters in our QFED model, we fix parameter $\lambda=0.05$ and adjust the others each time. $\alpha$ and $\beta$ are sequentially set values of [0.5, 1, 1.5, 2, 2.5, 3]. Then $\alpha$ and $\beta$ are set to 1.5 and 2 and the variation values of $\lambda$ are sequentially set values of [0.05, 0.06, 0.07, 0.08, 0.09, 0.1]. To effectively estimate the fusion performance, the combined quality metric $Q_{E}+Q_{CB}$ jointly assesses structural consistency and perceptual quality. Since $\lambda$ is not remarkably sensitive to the final fusion quality, $Q_{MI}$ is used to quantitatively analyze the fused results. We select the values corresponding to the maximum quality measure for the balance between detail preservation and background consistency.
Fig. 5 (a) displays the $Q_{E}+Q_{CB}$ results of QFED with the samples of mffw dataset. $\alpha$ and $\beta$ are set to 1.5 and 2 on mffw dataset as this setting yields the highest combined quality score. For lytro and MFI-WHU datasets, we select the optimal parameters $\alpha$ and $\beta$ in this manner and they are set to 1.5 and 0.5. In Fig. 5 (b), $Q_{MI}$ achieves its peak performance at $\lambda=0.05$. We adopt $\lambda=0.05$ consistently for all datasets since $\lambda$ exhibits less sensitivity to fusion quality variations. The parameter settings are uniformly employed in all subsequent experiments and evaluations.
TABLE II: Ablation study on each term of QFED. The best results are bolded.
Term settings $Q_{MI}\uparrow$ $Q_{G}\uparrow$ $Q_{P}\uparrow$ $Q_{E}\uparrow$ $Q_{CB}\uparrow$
$\bm{\alpha=0}$ 1.0580 0.7359 0.7488 0.8240 0.7298
$\bm{\beta=0}$ 1.0052 0.7338 0.7282 0.8279 0.6980
$\bm{\lambda=0}$ 1.0140 0.7324 0.7432 0.8259 0.7127
QFED 1.0685 0.7374 0.7832 0.8234 0.7508
Effectiveness of QFED. To verify the effectiveness of the QFED model, we perform the ablation study on each term of QFED. Specifically, three different settings of $\alpha=0$, $\beta=0$, $\lambda=0$ and $\alpha\neq 0$, $\beta\neq 0$, $\lambda\neq 0$ are particularly considered in this experiment. Table II gives the quantitative fusion results obtained by removing each term of QFED separately. As can be observed, each term can influence the fusion performance at different degrees. When $\beta$ is set to zero, we may obtain more focus-related features in the detail-scale layer $\dot{\bm{\mathrm{D}}}$. However, it fails to balance a tradeoff in the dual-scale focus map generation and subsequent fusion process. QFED provides the best quantitative results.
IV-C Performance Comparison
This subsection discusses the quantitative and qualitative comparison results of color image fusion.
Quantitative evaluation. Table III provides the average performance of various competing algorithms on three public datasets. Six evaluation metrics are employed to assess the quality of the fused images, with higher values indicating better performance. In this paper, the highest evaluation scores are highlighted in bold.
On the lytro dataset [35], our framework achieves the highest metric values compared to other competing methods across gradient-based, structural similarity-based, and human-perception-based evaluation metrics. This improvement is attributed to accurate focus detection capability of our QMCIF framework. However, our framework is worse than MCCSR-Net and SAMF on the metric of $Q_{MI}$. This is because our patch-wise fusion strategy tends to preserve less image intensity information during the quaternion fusion and refinement process compared to the pixel-wise fusion strategy.
[…]
@@ IV-C Performance Comparison
TABLE III: Performance comparison of multi-focus image fusion methods.
@@
SwinMFF [8] 0.7689 0.7013 0.7773 0.8257 0.8867 0.6413
MCCSR-Net [34] 1.1920 0.7619 0.8456 0.8772 0.9886 0.8084
Ours 1.1656 0.7603 0.8472 0.8814 0.9896 0.8108
mffw [29]
SR [14] 0.7181 0.6278 0.5500 0.8078 0.9349 0.6503
CSR [31] 0.9023 0.7065 0.6903 0.8222 0.8556 0.6861
@@
MFI-WHU dataset contains challenging small-area blurred structures. Our framework achieves the best fusion performance in $Q_{G}$ and $Q_{E}$ metrics, underscoring its effectiveness in detecting focus regions from small areas, particularly when compared to SAMF. Compared with QHOSVD and MCCSR-Net, QMCIF clearly preserves more accurate structural details and achieves the highest perceptual quality $Q_{CB}$. However, in the $Q_{Y}$ metric, our framework performs slightly worse than GFDF and MFIF-GAN since both methods exhibit strong pixel-wise focus detection capabilities and our patch-wise fusion strategy tends to preserve less information from the original image in the boundary regions.
@@
Extensive experiments on various datasets demonstrate that our framework significantly outperforms QHOSVD in terms of detail preservation and artifact removal under real-world challenging scenarios.
Visual evaluation. To assess the effectiveness of focus detection and artifact suppression, we compare our framework against five recent state-of-the-art fusion approaches. More visual comparison results are seen in the supplementary material.
The top row of Fig. 6 presents the fusion results on the lytro dataset. As observed in the red bounding boxes, QHOSVD and TCMOA suffer from noticeable edge blurring in the building structures while MCCSR-Net introduces spatial artifacts in the same regions. In contrast, our framework produces a visually pleasing fusion result with clear object boundaries and no observable artifacts.
The middle row of Fig. 6 shows the results on the mffw dataset. The blue bounding boxes reveal that QHOSVD, SAMF, and TCMOA generate blurred pixels near the object boundaries, and the DBMFIF output contains pseudo-edges and ghosting artifacts in low-contrast regions. Our framework delivers the most visually satisfactory fusion result, effectively addressing the challenges posed by low-gradient backgrounds and uncertain boundary transitions.
The bottom row of Fig. 6 displays fusion results on the MFI-WHU dataset. This scene contains small and near-blurred structures such as electric wires and poles, making focus estimation particularly challenging. The red bounding boxes clearly show that QHOSVD, SAMF, TCMOA, and MCCSR-Net fail to preserve these fine structures. In contrast, our framework accurately preserves these focused regions, demonstrating its superior capability in fine-grained focus detection and high-fidelity structure retention.
Our framework can be easily extended to multiple color image fusion with architectural modifications. As an example, we extend our QMCIF framework to fuse three partially focused color images. By converting the input images into multiple quaternion representations, we obtain the detail layer and coefficient matrix for each input quaternion representation individually and generate patch-wise dual-scale focus maps using Eqs. (16) and (17). We select the patch with the most focus level of input images. Then we obtain dual-scale fusion results and the final fused result simply applying the QBDF and QSSR strategies. More details of multiple color image fusion are seen in the supplementary material.
@@
Figure 7: Visual comparison of three partially focused color image fusion on lytro3 dataset.
Figure 8: Visual comparison of three partially focused color image fusion on mffw3 dataset.
We conduct experiments on the lytro3 [37] and mffw3 [29] datasets. We compare our framework with three representative approaches that are designed for multiple image fusion. The deep learning-based baseline ZMFF is used for this comparison [32]. Figs. 7 and 8 presents the visual comparison results. More comparison results are included in the supplementary material. In Fig. 7, QHOSVD produces visibly blurred shapes in the region marked by the red bounding box, especially around the device contour. ZMFF also suffers from color inconsistency in the same region, leading to degraded structure preservation and unnatural appearance. In Fig. 8, ZMFF exhibits noticeable visual artifacts in the background, degrading perceptual quality. In contrast, our framework produces clean and structurally consistent results across all scenes. It effectively eliminates visual artifacts and maintains accurate focus representation in both foreground and background regions, as well as at structural boundaries. Our framework offers a simple yet robust solution for complex fusion tasks.
IV-D Ablation study
In this subsection, we conduct an ablation study with different settings of our QMCIF framework on mffw dataset [29].
TABLE IV: Quantitative evaluation on different settings of our QMCIF framework.